OpenStudy (anonymous):

Find the product of z1 and z2 where z1 = 2(cos 80° + i sin 80°), and z2 = 9(cos 110° + i sin 110°). 11(cos 190° + i sin 190°) 18(cos 8800° + i sin 8800°) 18(cos 190° + i sin 190°) 18(cos 80° + i sin 80°)

jimthompson5910 (jim_thompson5910):

If z1 = r1*( cos(theta1) + i*sin(theta1) ) z2 = r2*( cos(theta2) + i*sin(theta2) ) then z1*z2 = (r1*r2) * ( cos(theta1 + theta2) + i*sin(theta1 + theta2) )

OpenStudy (anonymous):

18(cos 190° + i sin 190°) is this right?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

can u help me here? Express the complex number in trigonometric form. 6 - 6i six square root two times the quantity cosine of seven pi divided by four plus i times sine of seven pi divided by four six square root two times the quantity cosine of five pi divided by four plus i times sine of five pi divided by four six times the quantity cosine of seven pi divided by four plus i times sine of seven pi divided by four six times the quantity cosine of five pi divided by four plus i times sine of five pi divided by four

jimthompson5910 (jim_thompson5910):

6-6i is in the form a+bi a = 6 b = -6

jimthompson5910 (jim_thompson5910):

to convert to trigonometric form, you use these formulas

jimthompson5910 (jim_thompson5910):

r = sqrt(a^2 + b^2) theta = arctan(b/a)

OpenStudy (anonymous):

theta=arctan (-6/6) theta=pi/4 r=6 so it's either \[6(\cos \frac{ 7\pi }{ 4 } + i \sin \frac{ 7\pi }{ 4 }) \] \[6(\cos \frac{ 5\pi }{ 4 } + i \sin \frac{ 5\pi }{ 4 })\] am i right?

jimthompson5910 (jim_thompson5910):

(6,-6) is in quadrant IV, so you should have gotten -pi/4 which is coterminal to 2pi+(-pi/4) = 7pi/4

jimthompson5910 (jim_thompson5910):

so it's \[\large 6(\cos \frac{ 7\pi }{ 4 } + i \sin \frac{ 7\pi }{ 4 }) \]

OpenStudy (anonymous):

ok thanks